OpenStudy (loser66):

How much heat transfer to the skin by 25.0 g of steam onto the skin? The latent heat of vaporization of steam is L = 2.256*10^6 J/kg

OpenStudy (abb0t):

Can't you use Q = mc\(\Delta\)T?

OpenStudy (loser66):

yes, I have Q = mc delta t + mL I confused, not sure + or - mL

OpenStudy (abb0t):

condensation is exothermic. so +

OpenStudy (loser66):

the steam releases the heat, so the heat-loss will have - sign, right?

OpenStudy (abb0t):

wait a minute. I thought i was in the chemistry section. this is physics. Oops.

OpenStudy (abb0t):

but since the substance is giving up energy in the form of heat, the enthalpy change of the substance is negative.

OpenStudy (loser66):

let see : skin temperature : 34 m = 25 *10^-3 kg c =4190 J/kg.K \(\triangle T= -66) so, \[Q = 25*10^{-3}*4190*(-66)+25*10^{-3}*2.256*10^6 = 49486J\]

OpenStudy (loser66):

but they said I am wrong

OpenStudy (abb0t):

Why is it -66?

OpenStudy (loser66):

initial temperature is 100 , final is 34 so delta T = 34-100 =-66

OpenStudy (abb0t):

And did you keep your units consistent. Notice that you have kg and g's.

OpenStudy (loser66):

25g = 25 *10^(-3) kg

OpenStudy (abb0t):

2256(25) J = 56400 J 25 g of 100°C hot water to water at 34 ° C gives off Q = cm\(\Delta\)T c = 4.18 J per (g°K) Now can you finish it?

OpenStudy (loser66):

Thank you, let me try. I got exactly this number , but it said "incorrect" Ha!! just homework online. I can log out and ask my prof to get the reason.

OpenStudy (abb0t):

I don't have a calculator with me right now actually. I am doing this all on my flip phone. HAHA

OpenStudy (raffle_snaffle):



OpenStudy (raffle_snaffle):

Q=(0.025kg)(2.256x1-^6)